OpenStudy (anonymous):

Which statements are correct interpretations of this graph? Choose exactly two answers that are correct

OpenStudy (anonymous):

Helppp

OpenStudy (anonymous):

Basically which of the two problems are proportional

OpenStudy (anonymous):

a)\[\frac{ 2 }{ 3 }\] b)\[\frac{ 3 }{ 2 }\] c)\[\frac{ 1.5 }{ 2 }\] d) .7

OpenStudy (dtan5457):

See which of the statements are true. B looks correct, as every 2 tomatoes does indeed become 3 dietary fiber, based on the graph. that being said c must also be correct, as 3/2=1.5

OpenStudy (anonymous):

Agree anybody

OpenStudy (anonymous):

Yup b and c

OpenStudy (anonymous):

thxs to all